OpenStudy (anonymous):

If f(x) = 4x – 7 and g(x) = 2x + 4, evaluate f(x) + g(x) for x = –3.

OpenStudy (anonymous):

If x = -3, you simply substitute -3 in for anywhere you see an 'x'. f(-3) = 4(-3) - 7 = -12 - 7 = -19 g(-3) = 2(-3) + 4 = -6 + 4 = -2 f(-3) + g(-3) = -19 + (-2) = -21